# CURATRON: Complete Robust Preference Data for Robust Alignment of Large Language Models

### v1

@@ Abstract
This paper addresses the challenges of aligning large language models (LLMs) with human values via preference learning (PL), with a focus on the issues of incomplete and corrupted data in preference datasets. We propose a novel method for robustly and completely recalibrating values within these datasets to enhance LLMs resilience against the issues. In particular, we devise a guaranteed polynomial time ranking algorithm that robustifies several existing models, such as the classic Bradley–Terry–Luce (BTL) (Bradley and Terry,, 1952) model and certain generalizations of it. To the best of our knowledge, our present work is the first to propose an algorithm that provably recovers an $\epsilon$-optimal ranking with high probability while allowing as large as $O(n)$ perturbed pairwise comparison results per model response. Furthermore, we show robust recovery results in the partially observed setting. Our experiments confirm that our algorithms handle adversarial noise and unobserved comparisons well in both general and LLM preference dataset settings. This work contributes to the development and scaling of more reliable and ethically aligned AI models by equipping the dataset curation pipeline with the ability to handle missing and maliciously manipulated inputs.
@@ 1 Introduction
Large Language Models (LLMs) are highly advanced Artificial Intelligence (AI) systems capable of understanding, interpreting, and generating languages. The integration of AI chatbots like ChatGPT into our daily lives and businesses has had a profound impact on both society and industries (Eloundou et al.,, 2023). These models have evolved from being specialized tools in specific fields to versatile assets that are increasingly applied in everyday activities and diverse work environments (Nguyen et al.,, 2023). However, the success of GPTs/LLMs depends not only on their ability to generate responses and perform tasks well but also on their alignment with human values and expectations.
@@
Prior to ChatGPT, AI language models and chatbots often produced undesirable answers, making them unfit for public use. Examples of this problem include Microsoft’s Tay chatbot (Lee,, 2016), the National Eating Disorders Association’s Tessa (Mccarthy,, 2023), and Google’s Meena/LaMDA (Adiwardana et al.,, 2020). Although Meena/Lambda’s release was initially stopped due to not complying with Google’s AI safety and fairness guidelines, the project was later released publicly as Bard because of pressure from ChatGPT and Bing Chat (Mok,, 2023). Unfortunately, its factual error ended up costing Alphabet \$100 billion in market cap (Wittenstein,, 2023). The most recent AI incident at the time of writing this paper involved Gemini’s image generation AI model, which was found to be biased in favor of certain minority groups, leading Google to shut it down (Gordon,, 2024). Attempts have been made to address these problems by measuring the personality traits of large language models and modifying them through fine-tuning techniques (Karra et al.,, 2023). However, it was not until Instruct-GPT (Ouyang et al.,, 2022) was released with Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al.,, 2022), that marked a significant breakthrough in aligning LLMs effectively with human values. RLHF has paved the way for OpenAI’s ChatGPT to reshape industries, and GPT-4 to be considered a preliminary form of artificial general intelligence (AGI) (Bubeck et al.,, 2023), a type of AI that is vastly more intelligent than humans.
We show from the examples above that alignment is a crucial step in steering AI models toward the desired outcomes that align with human values/intentions. The prevalent method for aligning AI/LLMs currently involves preference learning (PL) through RLHF or Reinforcement Learning from AI Feedback (RLAIF) (Lee et al.,, 2023) using Proximal Policy Optimization (PPO) (Schulman et al.,, 2017), or alternatively, employing Direct Preference Optimization (DPO) (Rafailov et al.,, 2023). While PPO is a reinforcement learning (RL) technique within the RLHF pipeline, DPO directly integrates human preferences into the LLMs.
@@
PL is an effective method, but it comes with a cost. This method requires high-quality human preference data (Touvron et al., 2023b, ). Acquiring human preference for LLMs is a time-consuming and expensive process that requires lots of resources and hiring qualified human labelers to ensure data quality (Bai et al.,, 2022; Lee et al.,, 2023). For LLaMa-2, while there was very limited information in the development of the model, the training cost was estimated to be between \$20-45 million, with \$8 million spent solely on preference data (Lambert,, 2023). Even for large AI companies like Google, Meta, OpenAI, and Anthtropic, it is often not feasible to account for all human values and cultures. Recent research by OpenAI shows that RLHF may not be able to scale to supervise AGI models reliably (Burns et al.,, 2023). One solution for the scaling problem is to crowdsource preference data collection to the general public, which significantly reduces costs and provides a more diverse set of human values. However, scaling up the collection of human preferences using this type of crowdsourcing often results in poor-quality data that is incomplete or adversarially corrupted. Participants may unintentionally or intentionally provide incomplete, inaccurate, or harmful feedback, as highlighted by (Casper et al.,, 2023; Draws et al.,, 2021; Daniel et al.,, 2018; Cocos et al.,, 2017). Using such data directly to train AI can result in suboptimal AI systems that are biased, have misaligned behaviors, or can lead to outcomes that are unfair, discriminatory, or harmful (Figure 1). Deploying models trained with poor-quality preference data can erode trust in AI systems, especially if they are used in critical applications. Users may lose confidence in the model’s ability to perform tasks accurately and align with their expectations and ethical standards, as mentioned in the examples above. Therefore, ensuring the reliability and integrity of preference datasets for aligning strong LLMs and AGI becomes challenging as the process scales up with large-scale responses and unscreened participants.
@@
Approaching the issues, we consider the following learning problem. Suppose there are $n$ responses we wish to order based on a notion of comparison, between every pair of responses, with probabilistic outcomes. Further, we are given a set, =$\aleph=\{(i,j,\{y_{ij}^{k}\})\}$, consisting of $K$ independent pairwise comparison outcomes, denoted by $\{y_{ij}^{k}\}\in\{0,1\},k\in[K]$, between pairs of responses $(i,j)\subseteq[n]\times[n]$, a significant proportion of which might be corrupted by an adversary. In this passive learning setting, the concrete questions we wish to address are:
@@
    Is it possible to identify the pairs whose comparison results were corrupted by an adversary?
@@
    Further, does it work well in practice when we may also encounter unobserved data?
@@
Figure 1: CURATRON corrects incomplete and adversarially corrupted preference data to improve RLHF/DPO alignment results compared to using the raw initial preference data.
@@
    Algorithms & guarantees:
    • Under certain (information-theoretically tight) identifiability assumptions on the properties of the adversary, we develop a correct and efficient ranking method, Robust Preference Data for Reliably Alignment (RORATRON), that guarantees $\epsilon$-accurate high-probability learnability in a manner that is ‘robust’ and oblivious to the effects of the adversary. Our learning algorithm is provably characterized by polynomial time computational complexity.
    • In practice, it is often the case that not all pairs are compared, and even the observed pairwise comparison data could be adversarially corrupted – we also develop Complete Robust Preference Data for Reliable Alignment (CURATRON) and characterize the conditions for guaranteed robust recovery in this scenario. This results in a practical implication of enhancing preference data collection efficiency by automatically generating complete datasets from limited missing preference data.
@@
    Experiments: Finally, we support our theoretical results by showing robust ranking results on both synthetic and real-world experiments. Our experiments demonstrate the potential of our method in helping create large-scale AI/LLMs that are more accurately aligned with human values using minimal human effort as we achieved high reconstruction accuracy despite severe data missing and corruption.
@@ 2 Related Work
We now briefly present relevant work in: (1) LLM alignment with PL from human feedback, (2) ranking models and ranking algorithms that handle noise, and (3) robust subspace recovery methods, which will be needed for us to prove recovery results for ranking.
@@
PL was initially developed to train agents in simulated environments to perform nuanced behaviors that are hard-to-define but easy to observe and recognize (Christiano et al.,, 2017). It has recently been found successful in aligning LLMs to human intentions and values such as harmfulness, helpfulness, factuality, and safety. Some of the methods of PL in LLMs are RLHF (Ouyang et al.,, 2022), DPO (Rafailov et al.,, 2023), Sequence Likelihood Calibration with Human Feedback (SLiC-HF) (Zhao et al.,, 2023), Rejection Sampling Optimization (RSO) (Liu et al., 2024b, ), and Pairwise Cringe Optimization (PCO) and iterative DPO (iDPO) (Xu et al.,, 2023). Recent works also attempt to eliminate the need for pairwise human preferences. Kahneman-Tversky Optimization (KTO) (Ethayarajh et al.,, 2023) requires only binary feedback on LLM outputs. Listwise Preference Optimization (LiPO) (Liu et al., 2024a, ) treats model alignment as a listwise ranking problem.
However, these methods require high-quality human supervision through pairwise or ranking preference data, which is often noisy and not practical in real-world scenarios (Casper et al.,, 2023). Trying to address this issue, robust DPO (rDPO) (Chowdhury et al.,, 2024), conservative DPO (cDPO) (Mitchell,, 2023), and Identity Preference Optimization (IPO) (Azar et al.,, 2023) losses are proposed to solve the problem of noisy preference data. An alternative solution to reduce dependence on human data is to utilize Constitutional AI/AIF (Bai et al.,, 2022; Lee et al.,, 2023). This approach aims to minimize human inputs by using other LLMs/models (Jiang et al., 2023b, ), or by self-critiquing and improving its outputs (Chen et al.,, 2024; Yuan et al.,, 2024) based on user-defined principles.
Our proposed approach to scaling up PL is different from previous work. Instead of creating new loss functions or using other models/LLMs themselves to remove humans outside of the loop, we aim to make PL with human feedback easier to manage by explicitly solving the problems associated with it. We develop low-cost and computationally efficient algorithms that take into account the low-rank characteristic of pairwise human preference data, which helps us address issues such as incomplete and adversarial corruptions in human preference datasets.
@@
The Robust PCA (RPCA) problem (Netrapalli et al.,, 2014) addresses the following question: suppose we are given a data matrix $\mathbf{M}$ which is the sum of an unknown low-rank matrix $\mathbf{L}$ and an unknown sparse matrix $\mathbf{S}$, can we recover each of the component matrices? RPCA is a widely used technique due to its ability to handle low-rankness, which is commonly observed in various applications such as recommendation systems, visual analytics, and social networks (Vaswani et al.,, 2018). Recently, RPCA has also been applied to parameter-efficient fine-tuning in the LLM domain (Nikdan et al.,, 2024).
@@
3 Problem Setup and Solution Approach

### v2

# CURATRON: Complete and Robust Preference Data for Rigorous Alignment of Large Language Models

@@ Abstract
This paper addresses the challenges of aligning large language models (LLMs) with human values via preference learning (PL), focusing on incomplete and corrupted data in preference datasets. We propose a novel method for robustly and completely recalibrating values within these datasets to enhance LLMs’ resilience against the issues. In particular, we devise a guaranteed polynomial time ranking algorithm that robustifies several existing models, such as the classic Bradley–Terry–Luce (BTL) (Bradley and Terry,, 1952) model and certain generalizations of it. To the best of our knowledge, our present work is the first to propose an algorithm that provably recovers an $\epsilon$-optimal ranking with high probability while allowing as large as $O(n)$ perturbed pairwise comparison results per model response. Furthermore, we show robust recovery results in the partially observed setting. Our experiments confirm that our algorithms handle adversarial noise and unobserved comparisons well in both general and LLM preference dataset settings. This work contributes to the development and scaling of more reliable and ethically aligned AI models by equipping the dataset curation pipeline with the ability to handle missing and maliciously manipulated inputs.
@@ 1 Introduction
Large Language Models (LLMs) are highly advanced Artificial Intelligence (AI) systems capable of understanding, interpreting, and generating languages. The integration of AI chatbots like ChatGPT into our daily lives and businesses has profoundly impacted society and various industries (Eloundou et al.,, 2023; Brynjolfsson et al.,, 2023). These models have evolved from being specialized tools in specific fields to versatile assets that are increasingly applied in everyday activities and diverse work environments. However, the success of GPTs/LLMs depends not only on their ability to generate responses and perform tasks well but also on their alignment with human values and expectations.
@@
Early challenges in advanced LLMs management: Before ChatGPT, AI language models and chatbots often produced undesirable answers, making them unfit for public use. Examples of this problem include Microsoft’s Tay chatbot (Lee,, 2016), the National Eating Disorders Association’s Tessa (Mccarthy,, 2023), and Google’s Meena/LaMDA (Adiwardana et al.,, 2020). Although Meena/Lambda’s release was initially stopped due to not complying with Google’s AI safety and fairness guidelines, the project was later released publicly as Bard because of pressure from ChatGPT and Bing Chat (Mok,, 2023). Unfortunately, its factual error ended up costing Alphabet \$100 billion in market cap (Wittenstein,, 2023). Another recent AI incident at the time of writing this paper involved Gemini’s image generation AI model, which was found to be biased in favor of certain minority groups, leading Google to shut it down (Gordon,, 2024). Attempts have been made to address these problems by measuring the personality traits of large language models and modifying them through fine-tuning techniques (Karra et al.,, 2023). However, it was not until Instruct-GPT was released with Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al.,, 2022), that marked a significant breakthrough in aligning LLMs effectively with human values. RLHF has paved the way for OpenAI’s ChatGPT to reshape industries, and GPT-4 to be considered a preliminary form of artificial general intelligence (AGI) (Bubeck et al.,, 2023), a type of AI that is vastly more intelligent than humans.
Preference Learning (PL) as a solution: Our examples above demonstrate that managing AI has long been recognized as imperative within both businesses and research communities (Berente et al.,, 2021). However, both sectors have been struggling with the unpredictable and sometimes adverse behaviors that stem from the increased autonomy and learning capabilities of generative models. Fortunately, alignment through preference learning (PL) offers an effective way to guide these models toward outcomes that align with human values and intentions, much like the approach used to manage human agents. Current PL methods for AI and LLMs are largely achieved through fine-tuning with either human feedback (Ouyang et al.,, 2022) or AI feedback (Lee et al.,, 2023). This process typically utilizes reinforcement learning algorithms, such as Proximal Policy Optimization (PPO) (Schulman et al.,, 2017), as an indirect fine-tuning approach; or alternatively, a direct fine-tuning approach, such as Direct Preference Optimization (DPO) (Rafailov et al.,, 2023).
Bottlenecks and Risks of PL: PL has proven to be an effective approach in training large language models (LLMs), but it entails significant costs, particularly in terms of human preference data collection (Touvron et al., 2023b, ). The process of acquiring human preferences for LLMs is both time-intensive and costly, requiring substantial resources and the recruitment of qualified human labelers to ensure high-quality data (Bai et al.,, 2022; Lee et al.,, 2023). For instance, the training costs for LLaMa-2 were estimated at \$20-45 million, with approximately \$8 million dedicated solely to acquiring preference data (Lambert,, 2023). Even major AI companies, such as Google, Meta, OpenAI, and Anthropic, encounter challenges in aligning their models to reflect a wide range of human values and cultural contexts reliably. OpenAI has raised concerns that reinforcement learning with human feedback (RLHF) may not scale effectively to supervise superintelligent models (Burns et al.,, 2023), a problem that becomes even more daunting for smaller enterprises and individual researchers with limited resources.
One proposed solution to this scaling challenge is the use of automated evaluators, such as LLMs or agents acting as judges, to infer preferences. However, these models also inherit biases from the human data used to train them, making them prone to similar reliability issues and vulnerabilities to manipulation. Furthermore, there is ongoing debate about whether human involvement should be fully eliminated from the development loop.
Another alternative approach is crowdsourcing the data collection process of preference data, involving the general public to provide a more diverse set of human values at a lower expense (Anthropic, 2023a, ). While crowdsourcing has the potential to alleviate the financial burden, it again cannot solve the challenges related to data integrity. Managing the quality of the data collected from large, diverse groups can be difficult, often resulting in low-quality or intentionally manipulated data.
@@
From our discussion above, collecting high-quality human preferences is the bottleneck to building better AIs and LLMs. Participants may unintentionally or intentionally provide incomplete, inaccurate, or harmful feedback, as highlighted by (Casper et al.,, 2023; Draws et al.,, 2021; Daniel et al.,, 2018; Cocos et al.,, 2017). Using such data directly to train AI can result in suboptimal AI systems that are biased, have misaligned behaviors, or can lead to outcomes that are unfair, discriminatory, or harmful (Figure 1). Deploying models trained with poor-quality preference data can erode trust in AI systems, especially if they are used in critical applications. Users may lose confidence in the model’s ability to perform tasks accurately and align with their expectations and ethical standards.
Figure 1: CURATRON corrects incomplete and adversarially corrupted preference data to improve RLHF/DPO alignment results compared to raw initial preference data.
@@
Framing the problem: To address the challenges, we frame the following learning problem: Suppose there are $n$ distinct responses we wish to order based on pairwise comparisons with probabilistic outcomes, aimed at training AIs or large language models (LLMs) to learn human preferences. We are given a full set =$\aleph=\{(i,j,\{y_{ij}^{k}\})\}$ consisting of $K$ independent pairwise comparison outcomes, denoted by $\{y_{ij}^{k}\}\in\{0,1\},k\in[K]$. Between pairs of responses $(i,j)\subseteq[n]\times[n]$, a significant proportion of which might be corrupted by an adversary. In this passive learning setting, the concrete questions we wish to address are:
@@
    Is it possible to identify the comparisons whose results were corrupted by an adversary?
@@
    Further, does it work well in practice when we may also encounter unobserved or incomplete data?
@@
    Algorithms & guarantees: We propose the following modular algorithms that help in denoising preference data without interfering with the internal Supervised fine-tuning (SFT) or PL fine-tuning processes.
    • Under certain (information-theoretically tight) identifiability assumptions on the properties of the adversary, we develop a correct and efficient ranking method, Robust Preference Data for Rigorous Alignment (Algorithm 3 RORATRON), that guarantees $\epsilon$-accurate high-probability learnability in a manner that is ‘robust’ and oblivious to the effects of the adversary. Our learning algorithm is provably characterized by polynomial time computational complexity.
    • In practice, it is often the case that not all pairs are compared, and even the observed pairwise comparison data could be adversarially corrupted – we propose Complete Robust Preference Data for Rigorous Alignment (Algorithm 4 CURATRON) and characterize the conditions for guaranteed robust recovery in this scenario.
@@
    Experiments: Finally, we support our theoretical results by showing robust ranking results in practice. We also propose a data augmentation technique that further improves our results. Our experiments demonstrate the potential of our method in helping create large-scale AI/LLMs that are more accurately aligned with human values using minimal human effort as we achieved high reconstruction accuracy despite severe data missing and corruption.
@@ 2 Related Work
We briefly present relevant work in: (1) LLM alignment with PL from human feedback, (2) ranking models and ranking algorithms that handle noise, and (3) robust subspace recovery methods, which will be needed for us to prove recovery results for ranking.
@@
PL was initially developed to train agents in simulated environments to perform nuanced behaviors that are hard-to-define but easy to observe and recognize (Christiano et al.,, 2017). It has recently been found successful in aligning LLMs to human intentions and values such as harmfulness, helpfulness, factuality, and safety. Some of the methods of PL in LLMs are Reinforcement learning from human feedback (RLHF) or AI feedback (RLAIF) via PPO (Ouyang et al.,, 2022; Lee et al.,, 2023), Direct Preference Optimization (DPO) (Rafailov et al.,, 2023), Sequence Likelihood Calibration with Human Feedback (SLiC-HF) (Zhao et al.,, 2023), Rejection Sampling Optimization (RSO) (Liu et al., 2024b, ), and Pairwise Cringe Optimization (PCO) and iterative DPO (iDPO) (Xu et al., 2023c, ). Recent works also attempt to eliminate the need for pairwise human preferences. Kahneman-Tversky Optimization (KTO) (Ethayarajh et al.,, 2023) requires only binary feedback on LLM outputs. Listwise Preference Optimization (LiPO) (Liu et al., 2024a, ) treats model alignment as a listwise ranking problem.
However, these methods require high-quality human or AI supervision through pairwise or ranking preference data, which is often noisy and not practical in real-world scenarios (Casper et al.,, 2023). Trying to address this issue, robust DPO (rDPO) (Chowdhury et al.,, 2024), conservative DPO (cDPO) (Mitchell,, 2023), and Identity Preference Optimization (IPO) (Azar et al.,, 2023) losses are proposed to solve the problem of noisy preference data. An alternative solution to reduce dependence on human data is to utilize Constitutional AI/AIF (Bai et al.,, 2022; Lee et al.,, 2023). This approach aims to minimize human inputs by using other LLMs/models (Jiang et al., 2023b, ), or by self-critiquing and improving its outputs (Chen et al.,, 2024; Yuan et al.,, 2024) based on user-defined principles.
Our proposed approach to robustly scaling up PL is different from previous work. Instead of creating new loss functions or using other models/LLMs themselves to remove humans outside of the loop, we aim to make PL with human feedback easier to manage by explicitly solving the problems associated with it. We develop low-cost and computationally efficient algorithms that take into account the low-rank characteristic of pairwise human preference data, which helps us address the data issues of incomplete and adversarial corruptions.
@@
The Robust PCA (RPCA) problem (Netrapalli et al.,, 2014) addresses the following question: suppose we are given a data matrix $\mathbf{M}$ which is the sum of an unknown low-rank matrix $\mathbf{L}$ and an unknown sparse matrix $\mathbf{S}$, can we recover each of the component matrices? RPCA is a widely used technique due to its ability to handle low-rankness, which is commonly observed in various applications such as recommendation systems, visual analytics, and social networks (Vaswani et al.,, 2018). To deal with large-scale anomalies, Netflix developed Robust Anomaly Detection (RAD) for time series based on RPCA (Blog,, 2015). Recently, RPCA has also been applied to parameter-efficient fine-tuning in the LLM domain (Nikdan et al.,, 2024).
@@
3 Problem Setting and Solution Approach
@@ 3.2 Characterization of the Adversary
Claim 1.
[Upper bound on estimation error] Under Assumption 1 it is possible that $\operatorname{dist}(\widehat{\sigma},\mathbf{P^{c}})=O(1)$.
3.3 Solution Approach
Our proposed methods are designed to handle raw corrupted preference data, which may involve missing values or adversarial corruption. The goal is to recover data that closely resembles the original counterfactual data. This recovered data is then employed in subsequent fine-tuning steps such as DPO fine-tuning, as depicted in Figure 1. To accomplish this, we first map the raw preference data into the logit preference matrix subspace, leveraging its rank-2 structure as described by Rajkumar and Agarwal, (2016) (see subsection 3.1). In this transformed subspace, a matrix close to rank 2 indicates clean, healthy data, while higher ranks suggest the presence of noise or potential corruption. If potential issues are detected, we apply our algorithms to impute missing values and/or separate clean and noisy components. Once the processes are completed, the recovered data is converted to its original form for subsequent processing. This pipeline ensures a reliable foundation for training robust AI models, even when the initial preference data is significantly compromised. We now identify three scenarios/settings where missing and adversarially corrupted comparisons can affect the ranking of responses and tackle them in detail in subsequent sections.
  1. 1.
  2. 2.
    Partially observed and uncorrupted setting (Figure 2(b), Section 10): Not all pairs of responses are compared. This scenario can happen when data quality is prioritized over quantity in the data collection process. Observing all possible comparisons can be expensive and challenging, especially when there are many responses/items to compare as mentioned in Subsection 3.5. Algorithm 5 CORATRON (also called LRPR in (Rajkumar and Agarwal,, 2016)) can be used to solve this problem.
  3. 3.
    Partially observed and adversarially corrupted setting (Figure 2(c), Section 5): Both (1) not all pairs of responses are compared, and (2) some instances of the comparison results are adversarially corrupted. This scenario can happen when data quantity and quality are not met in the data collection process. This scenario will likely happen in a large crowd-sourced environment due to large-scale LLM responses and participants. Algorithm 4 CURATRON is proposed to solve this problem.
3.4 Alignment of LLMs using Direct Preference Optimization
We utilize Direct Preference Optimization (DPO) (Rafailov et al.,, 2023) to align LLMs in our experiments to examine the efficacy of our algorithms in solving incomplete and adversarially corrupted preference data problems. DPO is a widely used method for PL, due to its simplicity, stability, performance, and computational efficiency. DPO can be used for many objectives, such as preferences, values, and reasoning. Essentially, DPO is the direct fine-tuning method with the loss function that can be expressed as a mathematical equivalent to RLHF. Its purpose is to optimize preferences by classifying human preference data without resorting to reward modeling or reinforcement learning.
where $y_{w}$ and $y_{l}$ represent preferred and dispreferred responses, respectively. $\pi_{\text{ref}}$ is the reference policy, and $\pi_{\theta}$ is the new policy.
3.5 The Important of Response Diversity
Recent advances in prompt engineering, such as self-consistency (SC) (Wang et al., 2023b, ), tree-of-thoughts (ToT) (Yao et al.,, 2023; Long,, 2023), and graph-of-thoughts (GoT) (Besta et al.,, 2024), have demonstrated improved performance by encouraging language models to generate diverse sets of responses before arriving at the final answer. Havrilla et al., (2024) highlight the importance of dataset diversity in achieving output diversity when fine-tuning language models, eventually enhancing generalization capabilities.
Therefore, while sampling multiple answers per prompt is not mandatory when fine-tuning with offline DPO or its variations, doing so and accurately ranking these answers before fine-tuning are crucial, as they help models improve their performance and agility to better generalize to unseen scenarios. Once we have accurately ranked the diverse set of responses, we have more freedom to choose different data sampling strategies depending on the use case, including using all comparisons, best-of-n, or worst-of-n, among others.
(a) RORATRON: Fully observed and adversarially corrupted setting
(b) CORATRON: Partially observed and uncorrupted setting
(c) CURATRON: Partially observed and adversarially corrupted setting
Figure 2: Three algorithms to address scenarios where incomplete data and adversarial corruption can impact LLMs.
4 Fully Observed and Adversarially Corrupted Setting
4.1 Algorithm
  1. 1.
    RPCA (Procedure 1): Note that Step 3 of Algorithm 3 uses a matrix low-rank plus sparse decomposition subroutine. To obtain our recovery guarantee, it is sufficient to use the RPCA problem as a black-box method; for the precise details of this algorithm, we refer the reader to (Netrapalli et al.,, 2014). In particular, our analysis uses the noise-case guarantees in their paper. This is characterized by a (strongly-polynomial) running time of $O(n^{2}r^{2}\log(1/\epsilon))$ and guarantees $\epsilon$-recovery of the component matrices under the conditions of Assumption 1 and Lemma 2. Since $n$ is generally small and improving the dependence on n is not valuable in the regimes applicable in our setting, any alternative RPCA method can be used for Step 3.
  2. 2.
    $\gamma$-approximate pairwise ranking procedure (Procedure 2): Step 4 of Algorithm 3 calls a constant factor approximate ranking procedure. Specifically, we use the Copeland procedure (Copeland,, 1951) which has a $5$-approximation guarantee (Coppersmith et al.,, 2006) and involves sorting the responses according to a score of response $i$ given by $\sum_{j=1}^{n}\mathbbm{1}(\overline{P}_{ij}>1/2)$.